Given the following information about Elkridge Sporting Goods, Inc., construct a balance sheet for June 30, 2013. On that date the firm had cash and marketable securities of $25,135, accounts receivable of $43,758, inventory of $170,000, net fixed assets of $329,600, and other assets of $13,125. It had accounts payables of $67,855, notes payables of $36,454, long-term debt of $223,600, and common stock of $150,000. How much retained earnings did the firm have? Retained earnings??
Answer #1
Total assets = 25135+43758+170000+329600+13125 = $         5,81,618
Total liabilities = 67855+36454+223600 = $         3,27,909
Per the accounting equation
Total assets = Total liabilities + Equity
Substituting values we have
581618 = 327909+Equity
Equity = 581618-327909 = $         2,53,709
Equity = Common stock+Retained earnings
253709 = 150000+Retained earnings
Retained earnings = 253709-150000 = $         1,03,709